OpenStudy (anonymous):

could any one help me on my exam because its critical that i ace it for gradution

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

ok great

OpenStudy (anonymous):

but do you when is your exam and timing also

OpenStudy (anonymous):

well i can start it now its algera

OpenStudy (anonymous):

ok imma start with the first question then?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Evaluate: -12 + 6 -18 6 -72 -6

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

Solve for x: -2(x - 3) = 10 x = -2 x = 2 x = -8 x = 8

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

there was some connection problem sorry

OpenStudy (anonymous):

Greg combined different weights of fertilizer, soil, and compost to make 29 kilograms (kg) of potting soil for some plants. The weight of fertilizer was x kg. The weight of soil was 2 kg more than three times the weight of fertilizer, and the weight of compost was half the weight of fertilizer. This is represented in the equation below. x + (3x + 2) + x = 29 What was the difference in the weights of soil and compost Greg combined? 3 kg 6 kg 14 kg 17 kg

OpenStudy (anonymous):

its okay :)

OpenStudy (anonymous):

is the equation right?

OpenStudy (anonymous):

i formed equation and answer is 17

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which relation does not represent a function? (-4, 2), (1, 5), (-2, 0), (1, 6) (-5, 3), (-3, 1), (0, -4), (3, -6) (2, 0), (4, 6), (6, 8), (10, 6) (-3, 6), (-1, 1), (0, 1), (3, 7)

OpenStudy (anonymous):

hey i have only 1 hour left so do it quick

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh i think ans should be from the graphs just plot a rough graph and see if the lines cut y axis more than one time youll get the answer

OpenStudy (anonymous):

meanwhile gimme next

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 3x < -9 or x + 2 10 A number line with a closed circle on -3, shading to the left, and an open circle on 8, shading to the right. A number line with an open circle on -3, shading to the left, and a closed circle on 8, shading to the right. A number line with an open circle on -3, a closed circle on 8, and shading in between. A number line with a closed circle on -3, an open circle on 8, and shading in between

OpenStudy (anonymous):

second inequality is unclear

OpenStudy (anonymous):

2(x + y) = 2x + 2y is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (precal):

really? you don't know this one?????

OpenStudy (precal):

Please read the code of conduct, we are here to help you learn concepts. We are not here to help you cheat on a test.

OpenStudy (anonymous):

i said i needed help my to learn

OpenStudy (anonymous):

thank you precal

OpenStudy (precal):

really ok let's talk about all of these properties

OpenStudy (precal):

Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property what do they mean?

OpenStudy (anonymous):

actually i dont want your help

OpenStudy (precal):

really, I really want to help you learn

OpenStudy (anonymous):

no you dont understand what sitution im in you came off mean

OpenStudy (precal):

mean, really? how? what words show you that I am mean. I am showing my concern for your learning, for your future.

OpenStudy (anonymous):

i have a timed exam i need to finish i'll be the first from my family to gradute but it okay fine dont help

OpenStudy (precal):

commutative property states a+b=b+a or ab=ba

OpenStudy (anonymous):

its distributive property

OpenStudy (precal):

I was eventually going to state that one property x(y+z)=xy+xz

OpenStudy (anonymous):

@precal i respect that u show concern for others but we cant give the things way we want if the other person is not ready for it if he is true for himself hell get reward otherwise hell suffer

OpenStudy (precal):

@Meinme please read code of conduct and the goal of openstudy

OpenStudy (anonymous):

What is the slope between the two points (8, 15) and (3, 5)? -2 2

OpenStudy (anonymous):

okay is it there?

OpenStudy (anonymous):

i have not gone thru it!!! sry :(

OpenStudy (anonymous):

oh well thanks for you help so far..

OpenStudy (anonymous):

hey @ninahey is this a test you are giving or was asking help for assingment cuz dude here we are trying to help people with subjects they are weak in not to gain marks...........it wont help u rather worsen your situation later.........................i dont mean to e harsh but you yourself somwhere down there know it wont help

OpenStudy (anonymous):

help for my last assigmment

OpenStudy (precal):

@Meinme you are helping for a timed test

OpenStudy (precal):

i have a timed exam i need to finish i'll be the first from my family to gradute but it okay fine dont help quote from ninahey

OpenStudy (anonymous):

but he says its assignment!!! :0

OpenStudy (anonymous):

yes my last one